12x^3 +7x^2-5x=0
a. solve by factoring
b. what are the solutions(roots)

Answer by stanbon(75887)

12x^3 +7x^2-5x=0
a. solve by factoring
x(12x^2+7x-5) = 0
b. what are the solutions(roots)
x=0 is a root
12x^2+7x-5 = 0
x= [-7+-sqrt(49-4*12*-5)]/24
x= [-7+-sqrt289]/24
x= [[-7+-17]/24
x= -1 or x= 5/12
==================
Cheers,
Stan H.